OpenStudy (anonymous):

Equation: (x-4)/(x+2) = 7

myininaya (myininaya):

\[\frac{x-4}{x+2}*(x+2)=7(x+2)\] \[x-4=7x+14\] \[-4-14=7x-x\] -28=6x \[\frac{-28}{6}=\frac{6x}{6}\] \[x=\frac{-28}{6}=\frac{-14}{3}\]

OpenStudy (anonymous):

right so x = ?

myininaya (myininaya):

i don't understand the question? did i lose you somewhere?

OpenStudy (anonymous):

just wondering if x = -28 -14 ---- = ---- 6 3 Is the final answer?

myininaya (myininaya):

do you think is? i see a mistake up there -4-14 is not -28

myininaya (myininaya):

-4-14=7x-x -18=6x x=-18/6=-3

OpenStudy (anonymous):

so x = -3?

myininaya (myininaya):

do you see it says x=-18/6=-3?

OpenStudy (anonymous):

yep

myininaya (myininaya):

then why are you asking me what x= when i said x=-18/6=-3

myininaya (myininaya):

did you not understand how i got x?

OpenStudy (anonymous):

i get it now just wasn't sure.